OpenStudy (anonymous):

-(3x - 2) = -4(x + 1) - 4

OpenStudy (anonymous):

First thing to do is distribute the - and the -4 through the parenthesis

OpenStudy (anonymous):

can you do that ?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

im very bad at math

OpenStudy (anonymous):

take the - and multiply it through everything in the parenthesis -(3x - 2) = -3x + 6 -4(x + 1) = -4x - 4 Do you understand how I did this ?

OpenStudy (anonymous):

wait ..the first one is -3x + 2

OpenStudy (anonymous):

how can i find the answer.

OpenStudy (anonymous):

we are getting there.... -(3x - 2) = -4(x + 1) - 4 -3x + 2 = -4x - 4 - 4 combine like terms -3x + 2 = -4x - 8 add 4x to both sides -3x + 2 + 4x = -4x + 4x - 8 combine like terms x + 2 = -8 now subtract 2 from both sides x + 2 - 2 = -8 - 2 combine like terms x = -10 I did it step by step....do you have any questions ?

OpenStudy (anonymous):

no. thank you very very much. i couldnt figure it out

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

do you understand how I got the answer ?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."